Hello,
Please help me with the following. I am having a difficult time!

1. In the following exercise, find the coordinates of the vertex for the parabola defined by the given quadratic function.
f(x)=4x^2+8x+1
The vertex is? (Type an ordered pair)
Thanks!!

Answer by stanbon(75887)

In the following exercise, find the coordinates of the vertex for the parabola defined by the given quadratic function.
f(x)=4x^2+8x+1
-----------
The vertex occurs where x = -b/2a = -8/(2*4) = -1
---
f(-1) = 4(-1)^2 + 8(-1) + 1
f(-1) = 4-8+1
f(-1) = -3
-------------------
Vertex: (-1,-3)
======================
Cheers,
Stan H.